# Extrapolation for bilinear compact operators in the variable exponent setting

### v1

@@ 7.C. Bilinear $\omega$-Calderón–Zygmund operators
$T(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K(x,y_{1},\ldots,y_{m})f(y_{1})\ldots f(y_{m})\,\mathrm{d}y_{1}\ldots\mathrm{d}y_{m},$
@@
Notice that the modulus of continuity $\omega(t)=t^{\varepsilon}$ with $\varepsilon>0$ satisfies the Dini condition. This example appeared in the work of Grafakos and Torres [16], where they studied the (standard) Calderón–Zygmund operator. It is worth to mention that for the general $\omega$, the linear $\omega$-Calderón–Zygmund operator was studied by Yabuta in [35]. A further extension of this result was obtained by Maldonado and Naibo [28] in the bilinear case.
The authors of [5, Theorem 5.1] proved the compactness of commutators of m-linear $\omega$-Calderón–Zygmund operators on weighted Lebesgue spaces with constant exponents. We point out that one part of the proof of [5, Theorem 5.1] relied on showing the following unweighted compactness result.
7.1 Theorem ([5]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$, $1<p_{1},\ldots,p_{m}<\infty$ and $0<p<\infty$ with $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T,b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{p}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
@@
If $T$ is a bilinear $\omega$-Calderón–Zygmund operator, then for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ with $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$ and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$, it holds that $T$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is almost the same to the one given in [33, Theorem 2.8]. In particular, the main difference is that instead of the pointwise estimate of [33, Proposition 7.2] we make use of the following one (see [27, Theorem 6.1])
@@
We notice that one of the main ingredients of the proof of [33, Theorem 2.8] was the characterization of $\mathcal{A}_{\vec{p}(\cdot)}$ weights by the bilinear maximal operator $\mathcal{M}$ (see [33, Theorem 2.4]). According to [33, Remark 2.7], one can obtain the multilinear version of this latter result. We also refer to the work [6, Theorem 2.2], in which the characterization of the $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ weights was established with the aid of the multilinear fractional maximal operator $\mathcal{M}_{\alpha}$.
@@
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Let $t$ be a fixed constant with $t>1$. Let $j\in\{1,\ldots,m\}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ with
$\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$
and
$t<\min\{(p_{j})_{-}:~j=1,\ldots,m\},$
and for $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t}}$, it holds that $T,b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is similar to [6, Theorem 2.4]. Here, we provide the necessary changes. In particular, let $0<\delta<\varepsilon<1/m$. Then, by [27, page 95 and Theorem 7.1] we have the pointwise estimate
@@
where the last inequality follows by using Jensen’s inequality. The rest of the proof proceeds in a similar fashion as in [6, Theorem 2.4] with the difference that instead of [6, Theorem 2.3] we apply Theorem 7.2; we skip the details. ∎
Now, we are in position to apply our abstract main result Theorem 1.1 which gives immediately the following.
@@
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$ and $t$ be a fixed constant with $t>1$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$ with
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)}$
and
@@
Choose now constants $0<x_{1},x_{2}<1$ such that
$0<x_{1}+x_{2}<1.$
Then, Theorem 7.1 implies that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=\left(\frac{1}{x_{1}},\frac{1}{x_{2}}\right)$, $p_{1}(\cdot):=\frac{1}{x_{1}+x_{2}}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$.
Hence, by Theorem 1.1 we deduce the desired conclusion for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$. This implies immediately the desired conclusion also for $T,\vec{b}]_{\Sigma}$, since the sum of bilinear compact operators remains compact due to [2, Proposition 2]. ∎
We close this subsection by mentioning that there are several examples of the bilinear $\omega$-Calderón–Zygmund operators (see [5, Subsection 5.1]) which satisfy the assumptions and conclusion of Theorem 7.5.
@@ 7.D. Bilinear fractional Calderón–Zygmund operators
$|K_{\alpha}(x,y_{1},\ldots,y_{m})|\leq\frac{A}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha}},$
@@
    $|K_{\alpha}(x,y_{1},\ldots,y_{m})-K_{\alpha}(x^{\prime},y_{1},\ldots,y_{m})|\leq\frac{B|x-x^{\prime}|}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha+1}},$
@@
    $|K_{\alpha}(x,y_{1},\ldots,y_{j},\ldots,y_{m})-K_{\alpha}(x,y_{1},\ldots,y_{j}^{\prime},\ldots,y_{m})|\leq\frac{B|y_{j}-y_{j}^{\prime}|}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha+1}},$
@@
As noted in [1] in the special case $m=2$, given such a kernel $K_{\alpha}$, the $m$-linear operator $T_{\alpha}$ defined by
$T_{\alpha}(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K_{\alpha}(x,y_{1},\ldots,y_{m})f_{1}(y_{1})\ldots f_{m}(y_{m})\,\mathrm{d}y_{1}\ldots\mathrm{d}y_{m},\quad x\in\mathbb{R}^{n}$
@@
As observed in [1, page 3] in the special case $m=2$ (see also [5, Theorem 5.8]), an example of such an operator is the $m$-linear fractional integral operator $\mathcal{I}_{\alpha}$, also known as the $m$-linear Riesz potential, associated to the kernel
$K_{\alpha}(x,y_{1},\ldots,y_{m}):=\frac{1}{(|x-y_{1}|+\ldots+|x-y_{m}|)^{mn-\alpha}}.$
@@
It was shown in [19] that $\mathcal{I}_{\alpha}$ extends to a bounded $m$-linear operator $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})$ whenever $1<p_{1},\ldots,p_{m}<\infty$, $0<q<\infty$ and $\frac{\alpha}{n}=\frac{1}{p}-\frac{1}{q}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Thus, the same holds for arbitrary $T_{\alpha}$.
A result on the compactness of commutators of $T_{\alpha}$ on unweighted Lebesgue spaces in the bilinear case was first established in [1, Theorem 2.1]. Subsequently, [5, Theorem 5.8] extended [1, Theorem 2.1] to the general $m$-linear setting and (constant exponent) weighted Lebesgue spaces and explicitly allowed the target space to be only a quasi-Banach space. Let us note that while [5, Theorem 5.8] is formally stated only for $\mathcal{I}_{\alpha}$, the authors of [5] observe that the statement and their proof hold for arbitrary $T_{\alpha}$. Here we only need the unweighted version of [5, Theorem 5.8].
7.7 Theorem ([1, 5]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$. Let $1<p_{1},\ldots,p_{m}<\infty$ and let $0<q<\infty$ such that $\frac{1}{p}-\frac{1}{q}=\frac{\alpha}{n}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\ldots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{q}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
Our goal consists in generalizing Theorem 7.7 to weighted variable exponent Lebesgue spaces in the bilinear case. To achieve that, we first need a boundedness result over such spaces. For the special case of the $m$-Riesz potential, this is just [6, Theorem 2.4]. Our proof of the general case is only a slight adaptation of [6, Theorem 2.4].
@@
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Let $t$ be a fixed constant with $t>1$ and $\gamma:=\frac{\alpha}{n}<\frac{m}{t}$. Let $j\in\{1,\ldots,m\}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{m}$ and $q(\cdot)\in\mathrm{LH}\cap\mathscr{P}_{0}$ with
$\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\frac{\alpha}{n}$
@@
$t<\min\{(p_{j})_{-}:~j=1,\ldots,m\},$
and for $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is almost identical to [6, Theorem 2.4], so we explain only the necessary changes.
First of all, by [6, Theorem 2.3] we have that the $m$-linear Riesz potential $\mathcal{I}_{\alpha}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\ldots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly. Thus, by (7.6) we have that the same holds for arbitrary $T_{\alpha}$.
Let now $0<\delta<\varepsilon<\frac{n}{mn-\alpha}$. Then, by [36, Lemma 2.24] we have the pointwise estimate
@@
where the constant $C$ depends only on $n,m$, the constants of $T_{\alpha}$, $\varepsilon$ and $\delta$. We now pick $\varepsilon:=\frac{1}{m}$, which is clearly a permitted choice. The rest of the proof proceeds exactly as in [6, Theorem 2.4]; we omit the details. ∎
Applying now our abstract main result, Theorem 1.1, we deduce immediately the following.
@@
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$. Set $\gamma:=\frac{\alpha}{n}$. Let $t$ be a fixed constant with $1<t<\frac{2}{\gamma}$. Then, for all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$ and $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\frac{\alpha}{n}$
@@
$t<\min\{(p_{j})_{-}:~j=1,2\},$
and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b_{1}]_{e_{1}}$, $T_{\alpha},b_{2}]_{e_{2}}$ and $T_{\alpha},\vec{b}]_{\Sigma}$ map $L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
@@
Part (1) of the assumptions of Theorem 1.1 for $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$ is just the conclusion of Theorem 7.8 in the bilinear case (with an obvious change of notation).
Pick now constants $0<\widetilde{x}_{1},\widetilde{x}_{2}<1$ such that
$\frac{\alpha}{n}<\widetilde{x}_{1}+\widetilde{x}_{2}<1+\frac{\alpha}{n}.$
@@
Such a choice is possible, since $\frac{\alpha}{n}<2$. Then, Theorem 7.7 yields that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=\left(\frac{1}{\widetilde{x}_{1}},\frac{1}{\widetilde{x}_{2}}\right)$, $q_{1}(\cdot):=\frac{1}{\widetilde{x}_{1}+\widetilde{x}_{2}-\frac{\alpha}{n}}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$.

### v2

@@ 7.C. Bilinear $\omega$-Calderón–Zygmund operators
$T(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K(x,y_{1},\ldots,y_{m})f(y_{1})\cdots f(y_{m})\,\mathrm{d}y_{1}\cdots\mathrm{d}y_{m},$
@@
Notice that the modulus of continuity $\omega(t)=t^{\varepsilon}$ with $\varepsilon>0$ satisfies the Dini condition. This example appeared in the work of Grafakos and Torres [undefo], where they studied the (standard) Calderón–Zygmund operator. It is worth to mention that for the general $\omega$, the linear $\omega$-Calderón–Zygmund operator was studied by Yabuta in [undefal]. A further extension of this result was obtained by Maldonado and Naibo [undefac] in the bilinear case.
The authors of [undefd, Theorem 5.1] proved the compactness of commutators of m-linear $\omega$-Calderón–Zygmund operators on weighted Lebesgue spaces with constant exponents. We point out that one part of the proof of [undefd, Theorem 5.1] relied on showing the following unweighted compactness result.
7.1 Theorem ([undefd]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$, $1<p_{1},\ldots,p_{m}<\infty$ and $0<p<\infty$ with $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T,b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\cdots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{p}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
@@
If $T$ is a bilinear $\omega$-Calderón–Zygmund operator, then for all proper $2$-admissible quadruples $(\vec{p}(\cdot),\vec{p}(\cdot),\vec{1},\infty)$ and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$, it holds that $T$ maps $L^{p_{1}(\cdot)}(w_{1})\times\cdots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is almost the same to the one given in [undefai, Theorem 2.8]. In particular, the main difference is that instead of the pointwise estimate of [undefai, Proposition 7.2] we make use of the following one (see [undefab, Theorem 6.1])
@@
We notice that one of the main ingredients of the proof of [undefai, Theorem 2.8] was the characterization of $\mathcal{A}_{\vec{p}(\cdot)}$ weights by the bilinear maximal operator $\mathcal{M}$ (see [undefai, Theorem 2.4]). According to [undefai, Remark 2.7], one can obtain the multilinear version of this latter result. We also refer to the work [undefe, Theorem 2.2], in which the characterization of the $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ weights was established with the aid of the multilinear fractional maximal operator $\mathcal{M}_{\alpha}$.
@@
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Let $t$ be a fixed constant with $t>1$. Let $j\in\{1,\ldots,m\}$. Then, for all proper $m$-admissible quadruples $(\vec{p}(\cdot),\vec{p}(\cdot),\vec{1},\infty)$ with
$t<\min\{(p_{k})_{-}:~k=1,\ldots,m\},$
and for all $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t}}$, it holds that $T,b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\cdots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{p(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is similar to [undefe, Theorem 2.4]. Here, we provide the necessary changes. In particular, let $0<\delta<\varepsilon<1/m$. Then, by [undefab, page 95 and Theorem 7.1] we have the pointwise estimate
@@
where the last inequality follows by using Jensen’s inequality. The rest of the proof proceeds in a similar fashion as in [undefe, Theorem 2.4] with the difference that instead of [undefe, Theorem 2.3] we apply Theorem 7.2; we skip the details. ∎
Now, we are in position to apply our abstract main result Theorem 1.1 which gives the following.
@@
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$ and $t$ be a fixed constant with $t>1$. Then, for all proper $2$-admissible quadruples $(\vec{p}(\cdot),\vec{p}(\cdot),\vec{1},\infty)$ with
@@
Choose now a constant
$x_{1}>\max\{t,2\}$
Such a choice is always possible. Then, Theorem 7.1 implies that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=(x_{1},x_{1})$, $p_{1}(\cdot):=\frac{x_{1}}{2}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$.
Hence, by Theorem 1.1 we deduce the desired conclusion for each of the operators $T,b_{1}]_{e_{1}}$ and $T,b_{2}]_{e_{2}}$. This implies immediately the desired conclusion also for $T,\vec{b}]_{\Sigma}$, since the sum of bilinear compact operators remains compact due to [undefa, Proposition 2]. ∎
We close this subsection by mentioning that there are several examples of the bilinear $\omega$-Calderón–Zygmund operators (see [undefd, Subsection 5.1]) which satisfy the assumptions and conclusion of Theorem 7.5.
@@ 7.D. Bilinear fractional Calderón–Zygmund operators
$|K_{\alpha}(x,y_{1},\ldots,y_{m})|\leq\frac{A}{(|x-y_{1}|+\cdots+|x-y_{m}|)^{mn-\alpha}},$
@@
    $|K_{\alpha}(x,y_{1},\ldots,y_{m})-K_{\alpha}(x^{\prime},y_{1},\ldots,y_{m})|\leq\frac{B|x-x^{\prime}|}{(|x-y_{1}|+\cdots+|x-y_{m}|)^{mn-\alpha+1}},$
@@
    $|K_{\alpha}(x,y_{1},\ldots,y_{j},\ldots,y_{m})-K_{\alpha}(x,y_{1},\ldots,y_{j}^{\prime},\ldots,y_{m})|\leq\frac{B|y_{j}-y_{j}^{\prime}|}{(|x-y_{1}|+\cdots+|x-y_{m}|)^{mn-\alpha+1}},$
@@
As noted in [undef] in the special case $m=2$, given such a kernel $K_{\alpha}$, the $m$-linear operator $T_{\alpha}$ defined by
$T_{\alpha}(\vec{f})(x):=\int_{(\mathbb{R}^{n})^{m}}K_{\alpha}(x,y_{1},\ldots,y_{m})f_{1}(y_{1})\cdots f_{m}(y_{m})\,\mathrm{d}y_{1}\cdots\mathrm{d}y_{m},\quad x\in\mathbb{R}^{n}$
@@
As observed in [undef, page 3] in the special case $m=2$ (see also [undefd, Theorem 5.8]), an example of such an operator is the $m$-linear fractional integral operator $\mathcal{I}_{\alpha}$, also known as the $m$-linear Riesz potential, associated to the kernel
$K_{\alpha}(x,y_{1},\ldots,y_{m}):=\frac{1}{(|x-y_{1}|+\cdots+|x-y_{m}|)^{mn-\alpha}}.$
@@
It was shown in [undefs] that $\mathcal{I}_{\alpha}$ extends to a bounded $m$-linear operator $L^{p_{1}}(\mathbb{R}^{n})\times\cdots\times L^{p_{m}}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})$ whenever $1<p_{1},\ldots,p_{m}<\infty$, $0<q<\infty$ and $\frac{\alpha}{n}=\frac{1}{p}-\frac{1}{q}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Thus, the same holds for arbitrary $T_{\alpha}$.
A result on the compactness of commutators of $T_{\alpha}$ on unweighted Lebesgue spaces in the bilinear case was first established in [undef, Theorem 2.1]. Subsequently, [undefd, Theorem 5.8] extended [undef, Theorem 2.1] to the general $m$-linear setting and (constant exponent) weighted Lebesgue spaces and explicitly allowed the target space to be only a quasi-Banach space. Let us note that while [undefd, Theorem 5.8] is formally stated only for $\mathcal{I}_{\alpha}$, the authors of [undefd] observe that the statement and their proof hold for arbitrary $T_{\alpha}$. Here we only need the unweighted version of [undefd, Theorem 5.8].
7.7 Theorem ([undef, undefd]).
Let $b\in\mathrm{CMO}(\mathbb{R}^{n})$. Let $1<p_{1},\ldots,p_{m}<\infty$ and let $0<q<\infty$ such that $\frac{1}{p}-\frac{1}{q}=\frac{\alpha}{n}$, where $\frac{1}{p}:=\sum_{j=1}^{m}\frac{1}{p_{j}}$. Then, $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}}(\mathbb{R}^{n})\times\cdots\times L^{p_{m}}(\mathbb{R}^{n})$ into $L^{q}(\mathbb{R}^{n})$ compactly for every $j=1,\ldots,m$.
Our goal consists in generalizing Theorem 7.7 to weighted variable exponent Lebesgue spaces in the bilinear case. To achieve that, we first need a boundedness result over such spaces. For the special case of the $m$-Riesz potential, this is just [undefe, Theorem 2.4]. Our proof of the general case is only a slight adaptation of [undefe, Theorem 2.4].
@@
Let $b\in\mathrm{BMO}(\mathbb{R}^{n})$. Set $\gamma:=\frac{\alpha}{n}$. Let $t$ be a fixed constant with $1<t<\frac{m}{\gamma}$. Let $j\in\{1,\ldots,m\}$. Then, for all proper $2$-admissible quadruples $(\vec{p}(\cdot),q(\cdot),\vec{1},\infty)$ with
$t<\min\{(p_{k})_{-}:~k=1,\ldots,m\}$
@@
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
and for all $m$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b]_{e_{j}}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\cdots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly.
@@
The proof is almost identical to [undefe, Theorem 2.4], so we explain only the necessary changes.
First of all, by [undefe, Theorem 2.3] we have that the $m$-linear Riesz potential $\mathcal{I}_{\alpha}$ maps $L^{p_{1}(\cdot)}(w_{1})\times\cdots\times L^{p_{m}(\cdot)}(w_{m})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ boundedly. Thus, by (7.6) we have that the same holds for arbitrary $T_{\alpha}$.
Let now $0<\delta<\varepsilon<\frac{n}{mn-\alpha}$. Then, by [undefam, Lemma 2.24] we have the pointwise estimate
@@
where the constant $C$ depends only on $n,m$, the constants of $T_{\alpha}$, $\varepsilon$ and $\delta$. We now pick $\varepsilon:=\frac{1}{m}$, which is clearly a permitted choice. The rest of the proof proceeds exactly as in [undefe, Theorem 2.4]; we omit the details. ∎
Applying now our abstract main result, Theorem 1.1, we deduce the following.
@@
Let $\vec{b}=(b_{1},b_{2})$ be a pair of functions in $\mathrm{CMO}(\mathbb{R}^{n})$. Set $\gamma:=\frac{\alpha}{n}$. Let $t$ be a fixed constant with $1<t<\frac{2}{\gamma}$. Then, for all proper $2$-admissible quadruples $(\vec{p}(\cdot),q(\cdot),\vec{1},\infty)$ with
$q_{-}>1,$
$t<\min\{(p_{j})_{-}:~j=1,2\}$
@@
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma$
and for all pairs of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, it holds that $T_{\alpha},b_{1}]_{e_{1}}$, $T_{\alpha},b_{2}]_{e_{2}}$ and $T_{\alpha},\vec{b}]_{\Sigma}$ map $L^{p_{1}(\cdot)}(w_{1})\times L^{p_{2}(\cdot)}(w_{2})$ into $L^{q(\cdot)}(\nu_{\vec{w}})$ compactly.
@@
Part (1) of the assumptions of Theorem 1.1 for $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$ is just the conclusion of Theorem 7.8 in the bilinear case (with different notation).
Pick now a constant
$d_{1}\in\left(\max\left\{t,\frac{2}{1+\gamma}\right\},\frac{2}{\gamma}\right).$
@@
Such a choice is clearly possible. Then, Theorem 7.7 yields that part (2) of the assumptions of Theorem 1.1 is satisfied with the particular choices $\vec{p}_{1}(\cdot):=(d_{1},d_{1})$, $q_{1}(\cdot):=\frac{1}{\frac{2}{d_{1}}-\gamma}$ and $\vec{w}_{1}:=(1,1)$ for each of the operators $T_{\alpha},b_{1}]_{e_{1}}$ and $T_{\alpha},b_{2}]_{e_{2}}$.